Imagine that stock in Company XYZ is selling for $50 per share. There are a million shares of stock, and Company XYZ made $2 million in profits last year. What is its P/E ratio?

a. 1

b. 10

c. 20

d. 25

e. 50
Answer #1

Earnings per share = Net income / number of shares outstanding

Earnings per share = $2,000,000 / 1,000,000 = $2 per share

P / E ratio = Market price per share / Earnings per share

P/E ratio = $50 / $2

P/E ratio = 25 times